40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9382     Entry time: Thu Nov 14 02:50:43 2013     In reply to: 9378     Reply to this: 9401
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRM oplev measured and modeled TF 

In the process of figuring out what we can do to fix our PRM motion problem, I am looking at the PRM oplev. 

Eventually (as in, tomorrow), I'd like to be able to simulate some optic motion as a result of an impulse, and see what the oplev loops do to that motion.  (For starters, I'll take the impulse response of the OSEM loop as my time series that the oplev loop sees).

One thing that I have done is look at the oplev model that Rana put together, which is now in the noisebudget svn: /ligo/svncommon/NbSVN/aligonoisebudget/trunk/OpLev/C1

This script plots the open loop gain of the modeled oplev:

PRM_OL_TF_model.png

This should be compared to the pitch and yaw measured transfer functions:

 PRM_OLPIT_TF.pdf

PRM_OLYAW_TF.pdf

In the YAW plot, there are 2 transfer functions.  The first time around, the UGF was ~2.5Hz, which is too low, so I increased the gain in the C1:SUS-PRM_OLYAW filter bank from -3 to -9. 

The shapes of the measured and modeled transfer functions look reasonably similar, but I haven't done a plot overlay.  I suspect that the reason I don't see the same height peak as in the model is just that I'm not taking a huge number of points.  However, if the other parts of the TF line up, I'll assume that that's okay.

I want to make sure that the modeled transfer function matches the measured ones, so that I know I can trust the model.  Then, I'll figure out how to use the time series data with the simulated loop.  Ideally, I'd like to see that the oplev loop can fully squish the motion from the OSEM kicks.  Once I get something that looks good (by hand-tweaking the filter shape), I'll give it a try in the actual system.  We should, as soon as I get the optimal stuff working, redo this in a more optimal way.  Both now, and after I get an optimal design, I'll look at the actual step and impulse responses of the loop, to make sure there aren't any hidden instabilities.

Other thoughts for the night:

Rana suggests increasing the gain in some of the oplev QPD heads (including PRM), so that we're getting more than a few hundred counts of power on each quadrant.  Since our ADCs go to 32,000 counts, a few hundred is very small, and keeping us close to our noise limits.

Also, just an observation, but when I watch the REFL camera along with POP and AS, it's clear that the PRM is getting kicked, and I don't have the ETMs aligned right now, so this is just PRMI flashes.  There is also a lot of glow in the BS chamber during flashes (as seen on the PRM face video camera).

ELOG V3.1.3-